$SPAD/src/input kernel.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool kernel.output
)set message test on
)set message auto off
)clear all
--S 1 of 19
x :: Expression Integer
--R
--R
--R   (1)  x
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 19
kernel x
--R
--R
--R   (2)  x
--R                                            Type: Kernel(Expression(Integer))
--E 2
--S 3 of 19
sin(x) + cos(x)
--R
--R
--R   (3)  sin(x) + cos(x)
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 19
kernels %
--R
--R
--R   (4)  [sin(x),cos(x)]
--R                                      Type: List(Kernel(Expression(Integer)))
--E 4
--S 5 of 19
sin(x)**2 + sin(x) + cos(x)
--R

--R
--R              2
--R   (5)  sin(x)  + sin(x) + cos(x)
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 19
kernels %
--R
--R
--R   (6)  [sin(x),cos(x)]
--R                                      Type: List(Kernel(Expression(Integer)))
--E 6
--S 7 of 19
kernels(1 :: Expression Integer)
--R
--R
--R   (7)  []
--R                                      Type: List(Kernel(Expression(Integer)))
--E 7
--S 8 of 19
mainKernel(cos(x) + tan(x))
--R
--R
--R   (8)  tan(x)
--R                                 Type: Union(Kernel(Expression(Integer)),...)
--E 8
--S 9 of 19
height kernel x
--R
--R
--R   (9)  1
--R                                                        Type: PositiveInteger
--E 9
--S 10 of 19
height mainKernel(sin x)
--R
--R
--R   (10)  2
--R                                                        Type: PositiveInteger
--E 10
--S 11 of 19
height mainKernel(sin cos x)
--R
--R

--R   (11)  3
--R                                                        Type: PositiveInteger
--E 11
--S 12 of 19
height mainKernel(sin cos (tan x + sin x))
--R
--R
--R   (12)  4
--R                                                        Type: PositiveInteger
--E 12
--S 13 of 19
operator mainKernel(sin cos (tan x + sin x))
--R
--R
--R   (13)  sin
--R                                                          Type: BasicOperator
--E 13
--S 14 of 19
name mainKernel(sin cos (tan x + sin x))
--R
--R
--R   (14)  sin
--R                                                                 Type: Symbol
--E 14
--S 15 of 19
f := operator 'f
--R
--R
--R   (15)  f
--R                                                          Type: BasicOperator
--E 15
--S 16 of 19
e := f(x, y, 10)
--R
--R
--R   (16)  f(x,y,10)
--R                                                    Type: Expression(Integer)
--E 16
--S 17 of 19
is?(e, f)
--R
--R
--R   (17)  true
--R                                                                Type: Boolean

--E 17
--S 18 of 19
is?(e, 'f)
--R
--R
--R   (18)  true
--R                                                                Type: Boolean
--E 18
--S 19 of 19
argument mainKernel e
--R
--R
--R   (19)  [x,y,10]
--R                                              Type: List(Expression(Integer))
--E 19
)spool
)lisp (bye)

———-

References^{[1] nothing}

7